# Probabilistic estimates of the diameters of the Rubik’s Cube groups

### v3

@@ I Objective
The diameter of the Cayley graph of the Rubik’s CubeHofstadter (1981); Singmaster (1981); Ferenc (2024) group (hereafter the diameter) is the fewest number of turns to solve the Cube from any initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
@@ II Background
Rubik’s CubeHofstadter (1981); Singmaster (1981); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf,Korf (1983, 1985) especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theory.Zassenhaus (1982); Fiat et al. (1989); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) Furthermore, its similarities with problems appearing in quantum mechanics,Corli et al. (2021) statistical mechanics,Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014) high-energy physics,Czech et al. (2011) and cryptographyDiaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@ IV.1 Half-turn metric
With the front-top-left cubie held fixed, only the right ($R$), down ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation.Singmaster (1981)
@@ V.1 Half-turn metric
In the first step of the half-turn metric, the 3$\times$3$\times$3 Cube can undergo any one of the following eighteen (18) turns: clockwise 90 turns $R$, $D$, $B$, left ($L$), up ($U$), front ($F$), and their inverses $R^{-1}$, $D^{-1}$, $B^{-1}$, $L^{-1}$, $U^{-1}$, $F^{-1}$ as well as 180 turns $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, $F^{2}$ in the Singmaster notation.Singmaster (1981) In the second and subsequent steps, $X$, $X^{-1}$, and $X^{2}$ immediately following the same (where $X=R$, $D$, $B$, $L$, $U$, or $F$) only produce the configurations that are already visited. Furthermore, repeated commutative turns such as $RL$ and $LR$ lead to the same configurations, together indicating $r<15$.
@@ VI.1 Half-turn metric
There are twenty-seven (27) turns, which are clockwise 90 turns $R_{1}$, $D_{1}$, $B_{1}$, $R_{2}$, $D_{2}$, $B_{2}$, $R_{3}$, $D_{3}$, $B_{3}$ and their inverses $R_{1}^{-1}$, $D_{1}^{-1}$, $B_{1}^{-1}$, etc. as well as the corresponding 180 turns $R_{1}^{2}$, $D_{1}^{2}$, $B_{1}^{2}$, etc., where the subscript $i$ in $X_{i}$ refers to the $i$th layer rotated with $i=1$ being the exterior layer (and hence $R_{3}$ is alternatively denoted by $L_{2}^{-1}$).
@@ VII.1 Half-turn metric
The thirty-six (36) turns in this metric are exterior-layer clockwise 90 turns $R_{1}$, $D_{1}$, $B_{1}$, $L_{1}$, $U_{1}$, $F_{1}$, interior-layer clockwise 90 turns $R_{2}$, $D_{2}$, $B_{2}$, $L_{2}$, $U_{2}$, $F_{2}$, and their inverses as well as the corresponding 180 turns.
@@ VIII Discussion
Table 11 summarizes the size and metric dependence of the observed (i.e., correct) and predicted diameters of the $n\times n\times n$ Rubik’s Cube groups ($2\leq n\leq 5$). This table can be easily extended to greater $n$, different metrics, or other types of Cubes with essentially no additional computational cost.
Table 11: Observed and predicted diameters of the $n\times n\times n$ Rubik’s Cube group ($2\leq n\leq 5$) in various metrics.
@@
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Observed Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (101) (this work).
@@
VIII.1 Comparison with the exact results for the 2$\times$2$\times$2 Cube
Figure 1: The number of new (non-redundant) configurations of the 2$\times$2$\times$2 Cube generated in each step in the half-turn metric. A terminal plot (step 11 for the observed) corresponds to the diameter.
@@ VIII.1 Comparison with the exact results for the 2$\times$2$\times$2 Cube
In Figs. 14 are plotted the number of new, non-redundant configurations generated (“Observed”) in each step for the half-turn, quarter-turn, semi-quarter-turn, and bi-quarter-turn metrics, respectively, which are compared with the probabilistic estimates (“Predicted”). The “Observed” numbers were determined by a computer program.Hirata (2014) The “Predicted” numbers were obtained as
@@
In each of the metrics, the observed curve in earlier steps displays a geometrical growth $r^{t}$ with the step count $t$ and is accurately reproduced by the probabilistically predicted curve. The peaks of the curves tend to occur one step too early in the predicted curves. The subsequent rapid falloffs of the observed curves seem super-exponential with their slopes in the logarithm scale becoming more negative for higher step counts. It may be imagined that in the final few steps of the diameter, the remaining unrealized configurations will be more efficiently exhausted than a random generation can. However, the predicted curves, owing to the non-Markovian assumption of the probabilistic estimation, will display an exponential and thus slower decay of the form $e^{-rt}$, where $r$ is the branching ratio, taking more steps than reality to exhaust the remaining unrealized configurations.
@@
In the half-turn and quarter-turn metrics, therefore, there seems to be a systematic cancellation of errors between the peak positions (occurring one step too early in the predicted curves) and the rate of decay (which is faster in the observed curves), although the predicted diameter in the half-turn metric is still overestimated by one. In the semi-quarter-turn metric, since $r$ is small, there is a longer decay phase, causing the slow decay of the predicted curve to overshoot the diameter by as much as two, despite its peak position occurring one step too early. In the bi-quarter-turn metric, in contrast, $r$ is large, which minimizes the error of the predicted decay rate, and underestimates the diameter by one.
@@
Implicit in our probabilistic estimation logic is the assumption that the details of each metric are unimportant and all of its characteristics can be encapsulated into the branching ratio $r$, which is invariant of the step count. From Figs. 14, this assumption seems valid for the growth phase, but is no longer so in the decay phase. The fact that these curves show a super-exponential decay suggests that the effective value of $r$ increases near the diameter and may even exceed the number of turns. In fact, the observed and predicted curves in Fig. 3 behave like $e^{-4.0t}$ and $e^{-2.6t}$ at the diameter, respectively, and the factor of 4.0 in the former exceeds $r=2.77$ or even the number of turns, which is three. There seems room for improvement in the probabilistic estimates in the peak position and decay rate of these curves.
@@
In the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes, whose branching ratio $r$ is small relative to the huge $E[T_{N}]$, there are long decay phases, potentially causing considerable overestimation of the diameters.
@@ Acknowledgements.
The author thanks Mr. Suchen He (Peking University) and Professor P. Sadayappan (University of Utah) for discussions. The author is a Guggenheim Fellow of the John Simon Guggenheim Memorial Foundation.

### v4

@@ I Objective
The diameter of the Cayley graph of the Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) group (hereafter the diameter) is the fewest number of turns to solve the Cube from any initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
@@ II Background
Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf,Korf (1983, 1985) especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theory.Zassenhaus (1982); Fiat et al. (1989); Joyner (2008); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) Furthermore, its similarities with problems appearing in quantum mechanics,Corli et al. (2021) statistical mechanics,Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014) high-energy physics,Czech et al. (2011) and cryptographyDiaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@ IV.1 Half-turn metric
With the front-top-left cubie held fixed, only the right ($R$), downward ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation.Singmaster (1981)
@@ V.1 Half-turn metric
In the first step of the half-turn metric, the 3$\times$3$\times$3 Cube can undergo any one of the following eighteen (18) turns: clockwise 90 face layer turns $R$, $D$, $B$, left ($L$), upward ($U$), front ($F$), and their inverses $R^{-1}$, $D^{-1}$, $B^{-1}$, $L^{-1}$, $U^{-1}$, $F^{-1}$ as well as 180 turns $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, $F^{2}$ in the Singmaster notation.Singmaster (1981) In the second and subsequent steps, $X$, $X^{-1}$, and $X^{2}$ immediately following the same (where $X=R$, $D$, $B$, $L$, $U$, or $F$) only produce the configurations that are already visited. Furthermore, repeated commutative turns such as $RL$ and $LR$ lead to the same configurations, together indicating $r<15$.
@@ VI.1 Half-turn metric
There are twenty-seven (27) turns, which are clockwise 90 turns $R_{1}$, $D_{1}$, $B_{1}$, $R_{2}$, $D_{2}$, $B_{2}$, $R_{3}$, $D_{3}$, $B_{3}$ and their inverses $R_{1}^{-1}$, $D_{1}^{-1}$, $B_{1}^{-1}$, etc. as well as the corresponding 180 turns $R_{1}^{2}$, $D_{1}^{2}$, $B_{1}^{2}$, etc., where the subscript $i$ in $X_{i}$ refers to the $i$th layer rotated with $i=1$ being the face (exterior) layer (and hence $R_{3}$ is alternatively denoted by $L_{2}^{-1}$).
@@ VII.1 Half-turn metric
The thirty-six (36) turns in this metric are face (exterior) layer clockwise 90 turns $R_{1}$, $D_{1}$, $B_{1}$, $L_{1}$, $U_{1}$, $F_{1}$, interior layer clockwise 90 turns $R_{2}$, $D_{2}$, $B_{2}$, $L_{2}$, $U_{2}$, $F_{2}$, and their inverses as well as the corresponding 180 turns.
@@ VIII Discussion
Table 11 summarizes the size and metric dependence of the actual (if available) and predicted diameters of the $n\times n\times n$ Rubik’s Cube groups ($2\leq n\leq 5$). This table’s predictions can be easily extended to greater $n$, different metrics, or other types of Cubes with essentially little additional computational cost, but determining actual values will be formidable.
Table 11: Actual and predicted diameters of the $n\times n\times n$ Rubik’s Cube group ($2\leq n\leq 5$) in various metrics.
@@
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Actual Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (101) (this work).
@@
VIII.1 Comparison between predicted and actual results
Figure 1: The number of new (non-redundant) configurations of the 2$\times$2$\times$2 Cube generated in each step in the half-turn metric. A terminal plot (step 11 for the actual) corresponds to the diameter.
@@ VIII.1 Comparison between predicted and actual results
In Figs. 14 are plotted the number of new, non-redundant configurations generated (“Actual”) in each step for the half-turn, quarter-turn, semi-quarter-turn, and bi-quarter-turn metrics, respectively, which are compared with the probabilistic estimates (“Predicted”). The “Actual” numbers were determined by a computer program.Hirata (2014) The “Predicted” numbers were obtained as
@@
In each of the metrics, the actual curve in earlier steps displays a geometrical growth $r^{t}$ with the step count $t$ and is accurately reproduced by the probabilistically predicted curve. The peaks of the curves tend to occur one step too early in the predicted curves. The subsequent rapid falloffs of the actual curves seem super-exponential with their slopes in the logarithm scale becoming more negative for higher step counts. It may be imagined that in the final few steps of the diameter, the remaining unrealized configurations will be more efficiently exhausted than a random generation can. However, the predicted curves, owing to the non-Markovian assumption of the probabilistic estimation, will display an exponential and thus slower decay of the form $e^{-rt}$, where $r$ is the branching ratio, taking more steps than reality to exhaust the remaining unrealized configurations.
@@
In the half-turn and quarter-turn metrics, therefore, there seems to be a systematic cancellation of errors between the peak positions (occurring one step too early in the predicted curves) and the rate of decay (which is faster in the actual curves), although the predicted diameter in the half-turn metric is still overestimated by one. In the semi-quarter-turn metric, since $r$ is small, there is a longer decay phase, causing the slow decay of the predicted curve to overshoot the diameter by as much as two, despite its peak position occurring one step too early. In the bi-quarter-turn metric, in contrast, $r$ is large, which minimizes the error of the predicted decay rate, and underestimates the diameter by one.
@@
Implicit in our probabilistic estimation logic is the assumption that the details of each metric are unimportant and all of its characteristics can be encapsulated into the branching ratio $r$, which is invariant of the step count. From Figs. 14, this assumption seems valid for the growth phase, but is no longer so in the decay phase. The fact that these curves show a super-exponential decay suggests that the effective value of $r$ increases near the diameter and may even exceed the number of turns. In fact, the actual and predicted curves in Fig. 3 behave like $e^{-4.0t}$ and $e^{-2.6t}$(where $t$ is the number of steps) at the diameter, respectively, and the factor of 4.0 in the former exceeds $r=2.77$ or even the number of turns, which is three. There seems room for improvement in the probabilistic estimates in the peak position and decay rate of these curves.
Figure 5: The number of new (non-redundant) configurations of the 3$\times$3$\times$3 Cube generated in each step in the half-turn metric. The actual values (those in the last five steps were approximate) were taken from Table 5.1 of Rokicki et al.Rokicki et al. (2013)
For the 3$\times$3$\times$3 Cube, Rokicki et al.Rokicki et al. (2013) gave the actual and estimated numbers (“positions”) of new configurations generated at each step. They are compared with our predicted values in Fig. 5. In this case, the rising part and the peak of the actual curve is accurately reproduced by the prediction, but the falling part decays too slow in the latter, overestimating the diameter by two, although the Rokicki et al.’s values in the last five steps are approximate. The actual curve has a rapid falloff of $e^{-22.3t}$, although it likely decreases faster than exponentially. On the other hand, the predicted curve decays exponentially as $e^{-13.3t}$ corresponding to $r=13.33$. It again suggests that our probabilistic logic tends to underestimate the accerelated rate of exhausting remaining unrealized configurations in the final few steps. However, the exact agreement of the predicted and actual diameters in the quarter-turn metric may contradict this view, given the smaller value of $r=9.37$ and even slower decay of the predicted curve in this metric.
@@
In the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes, whose branching ratio $r$ is small relative to the huge $E[T_{N}]$, there expected to be long decay phases, potentially causing noticeable overestimation of the diameters.
@@ Acknowledgements.
The author is a Guggenheim Fellow of the John Simon Guggenheim Memorial Foundation.